# The Illusive Slump of Disruptive Patents

### v1

@@
Abstract
Despite tremendous growth in the volume of new scientific and technological knowledge, the popular press has recently raised concerns that disruptive innovative activity is slowing. These dire prognoses were mainly driven by Park et al. (2023), a Nature publication that uses decades of data and millions of observations coupled with a novel quantitative metric (the CD index) that characterizes innovation in science and technology as either consolidating or disruptive. We challenge the Park et al. (2023) methodology and findings, principally around concerns of truncation bias and exclusion bias. We show that 88 percent of the decrease in disruptive patents over 1980-2010 reported by the authors can be explained by their truncation of all backward citations before 1976. We also show that this truncation bias varies by technology class. We update the analysis to 2016 and account for a change in U.S. patent law that allows for citations to patent applications in addition to patent grants, which is ignored by the authors in their analysis. We show that the number of highly disruptive patents has increased since 1980—particularly in IT technologies. Our results suggest caution in using the Park et al. (2023) methodology as a basis for research and decision making in public policy, industry restructuring or firm reorganization aimed at altering the current innovation landscape.
@@ 1 Introduction
Continued innovation in science and technology is considered a bedrock for and driving force of growth and prosperity in most economies. A paper co-authored by Park, Leahey and Funk entitled, “Papers and Patents are Becoming Less Disruptive Over Time” was recently published in Nature (2023). Given its provocative title and findings, as well as the topic examined, the paper attracted significant and global media attention. Park et al. (2023) has been featured in hundreds of international newspapers and magazines Altmetric (2023): for example, The Economist (2023) emphasized in a report on the changing nature of science that "Papers and Patents are Becoming Less Disruptive", indicating in a subtitle that "why that is, is a mystery"; The New York Times (2023) further stated: "What Happened to All of Science’s Big Breakthroughs?"; and the Financial Times (2023) noted: "Science is Losing its Ability to Disrupt".
@@
One reason for this attention is that the Park et al. (2023) results show marked declines in disruptive innovation over time. The authors state in the abstract that their "results suggest that slowing rates of disruption may reflect a fundamental shift in the nature of science and technology" (p. 138), and note in the conclusion that "this trend is unlikely to be driven by changes in citation practices or the quality of published work. Rather, the decline represents a substantive shift in science and technology, one that reinforces concerns about slowing innovative activity. We attribute this trend in part to scientists’ and inventors’ reliance on a narrower set of existing knowledge" (p. 142).
@@
Another reason for this attention is that the Park et al. (2023) results have implications regarding the organization of the entire science and technology innovation process—from government research labs and universities to private and public enterprises. The authors’ methodology and findings unsurprisingly piqued the interests of researchers, commentators, and journalists on a global scale who offered myriad explanations, such as increased pressures on scientists to apply for large (interdisciplinary) projects, rising administrative burdens, declining basic research funding, increasing risk-aversion among scientists, and pressures to publish rapidly Rust (2023); Yanai and Lercher (2023). Still other researchers raised concerns around the Park et al. (2023) methodology as a measure of disruptive innovation Ruan et al. (2021); Bornmann et al. (2020a, b).
@@
An important question arises, however, as to whether the argument made by Park et al. (2023) is accurate. We have major concerns around their methodological approach regarding truncation and exclusion bias related to measurement. We show using patent data that the worrisome result suggested by the authors is instead mainly a consequence of the omission of citations to older innovations: i.e., the authors artificially truncate all backward citations of patents before 1976. We also show using patent data that substantive differences in the results occur from not considering patent law changes: i.e., the authors neglect backward citations of patents published after 2000 to patent applications. Both of these biases have large measurement effects on their proposed CD index and direct implications on the accuracy of their findings and conclusions.
In what follows, we first compare the Park et al. (2023) methodology with truncation to our own methodology with no truncation. We explain why truncation bias can have such a significant impact, and then assess the effects of truncation bias using patent data. We then update the patent data over 2011-2016 and take into account a change in patent law that allows for citations to patent applications as well as to patent grants. We show that the number of highly disruptive patents has increased since 2005. We conclude with a discussion of our main findings and the dangers of measurement bias in influencing public policy, industry organization, and firm organization decisions and interventions.
@@ 2 Measuring Patent Disruptiveness
The CD index was developed by Funk and Owen-Smith (2017) and used by Park et al. (2023) to characterize whether a patent or scientific publication is considered more consolidating (i.e., building upon previous research and reinforcing the status quo) or more disruptive (i.e., obsolescing previous research and pushing into new directions). [^1] Since its introduction by Funk and Owen-Smith (2017) the index has been used in a wide variety of analyses to capture the disruptive content of patents (e.g. Kaltenberg et al., 2023; Wu et al., 2019) and of scientific papers (e.g. Bornmann et al., 2020a; Wu et al., 2019). It is based on the idea that a patent or scientific publication is disruptive if "the subsequent work that cites it is less likely to also cite its predecessors" (Park et al., 2023, p. 139). The CD index ranges from -1 (consolidating) to 1 (disruptive). The measure uses five-year post-publication windows in its construction, referred to as $CD_{5}$. For scientific papers, it starts in 1945; for patents, it starts in 1980.
Fig. 1 plots two average $CD_{5}$ indices for approximately $3.66$ million patents over $1980$-$2010$ from the United States Patent and Trademark Office (USPTO): one based on the Park et al. (2023) methodology; the other based on our methodology. [^2] As in Park et al. (2023), we consider only utility patents and draw our data from PatentsView (version February $21$, $2023$). The database contains all US patents that have been published between $1976$ and September $29$, $2022$. While Park et al. (2023) focus on patents of the aggregate NBER technology fields ”Chemical”, ”Computer and Communications”, ”Pharmaceutical and Medical”, ”Electrical and Electronic”, and ”Mechanical”, which encompasses a total of $3,046,672$ granted utility patents over 1980-2010, our main analysis utilizes all granted utility patents published by the USPTO during the same period, resulting in a total of $3,662,051$ patents. For more details, see Appendix A. As is readily apparent, the evolution of each index differs markedly: the Park et al. (2023) methodology produces an average annual CD index that starts at 0.39 in 1980 (the first observation year) but declines rapidly over time; our methodology instead indicates an average annual CD index that starts at 0.09 in 1980 and declines more gradually over time. Fig. 1 nonetheless shows marked convergence in the two indices in the later years of the sample—particularly since 2000.
Figure 1: The CD index: Park et al. (2023) Methodology Versus Our Methodology
@@
The difference in the average CD indices arises from Park et al. (2023) truncating all backward citations before 1976. To the best of our understanding, the authors do not explain or acknowledge this truncation explicitly despite it having important consequences. The Park et al. (2023) truncated CD index has much larger values compared to our untruncated CD index for those patents published closer to the truncation year. The reason is simple: patents published closer to the truncation year (e.g., 1980) frequently cite patents published in proximate years (e.g., pre-1976), while patents published well-past the truncation (e.g., 2010) infrequently cite those patents. In the earlier years, many backward citations are truncated; in the later years, only a few are. As a result, the CD indices diverge in the early years and converge in the latter years.
This implies that the Park et al. (2023) results are significantly biased due to the truncation of backward patent citations, with the bias strongest for those patents published closer to the truncation year. Hence, their findings of a sharp decline in the average patent disruptiveness can mostly be attributed to a reduction in measurement bias; not to a decline in disruptive innovations. After correcting for truncation bias, our methodology shows that the CD index has a much lower starting point and more modest decline over time. Specifically, 88 percent of the total decline in the average disruptiveness of patents found by Park et al. (2023) can be explained by the authors artificially truncating all backward citations to patents published before 1976.
@@ 3 Why Truncation Matters
An example helps illustrate our argument: We compare two patents that both appear in the Park et al. (2023) dataset: [^3] Fig. B.3 in the Appendix shows the first page from each patent document listing the backward citations. US 4181011 was issued in 1980; US 6511791 was issued in 2003. US 4181011 makes 13 citations to patents published between 1958 and 1974; US 6511791 makes 11 citations to patents published between 1987 and 2001. Using the Park et al. (2023) methodology and replication data ($patentsview\_analytical\_df.csv$), the number of backward citations for patent US 4181011 is zero—given the exclusion of backward citations to patents published before 1976. In contrast, the number of backward citations for patent US 6511791 is 11—the same number as listed in the patent document.
Figure 2: How Truncation of Backward Citations Affects the CD Index
@@
Formally, this can be shown as follows. Park et al. (2023) define the CD index for a focal patent as:
$CD_{t}=\frac{1}{N}\sum_{N=i}^{N}(-2f_{it}b_{it}+f_{it}),\leavevmode\nobreak\ with$ (1)
$f_{it}=\begin{cases}1,&\text{if $i$ cites the focal patent within $t$ years ofpost-publication of the focal patent}\\ 0,&\text{otherwise}\end{cases}$
@@
$b_{it}=\begin{cases}1,&\parbox[t]{346.89731pt}{if $i$ cites parts of the predecessors of the focal patent within $t$ years of post-publication of the focal patent}\\ 0,&\text{otherwise}\end{cases}$
where $N$ is the sum of the number of patents citing: the focal patent only ($N_{F}$), the focal patent and parts of its predecessors ($N_{B})$, and parts of its predecessors only ($N_{R}$) within $t$ years after the focal patent publication. A forward citation increases the consolidating (disruptive) nature of the focal patent if the forward cited patent does (does not) cite the predecessor patents (i.e., those patents also cited by the focal patent).
We extend the CD index by directly considering truncation. Truncation affects the CD index in two ways. First, left-truncation (i.e., with respect to the earlier years) reduces the number of counted forward citations $N$ if the truncated backward citations are cited within the $t$ post-publication years of the focal patent:
$N^{tr}<N^{non-tr},$ (2)
where $tr$ stands for truncated and $non-tr$ for non-truncated backward citations.
Second, left-truncation reduces the $b_{it}$ value if a forward citation of the truncated predecessor is also a forward citation of the focal patent within $t$ years of publication. In this case:
@@
Considering equations (2) and (3), we distinguish among four distinct cases: (i) no backward citations are truncated; (ii) backward citations are truncated but none are cited within the $t$ post-publication years of the focal patent; (iii) truncated backward citations are cited within the $t$ post-publication years of the focal patent but none cite the focal patent; and (iv) the truncated backward citations are cited within the $t$ post-publication years of the focal patent and at least one of them cite the focal patent. In case (i) and case (ii), equations (2) and (3) hold as equalities and truncation has no effect on the CD index:
@@
In case (iii), the inequality of equation (2) holds but equation (3) is strictly an equality. In case (iv), the inequalities of equations (2) and (3) hold. In case (iii) and case (iv), the truncated CD index is biased upward as long as the summation part of equation (1) of the truncated CD index is positive, because each of the two effects artificially increases its value. Moreover, ceteris partibus, the bias is stronger in case (iv). Both results are easily seen by adding the inequalities of equations (2) and (3) to the CD index provided in equation (1). As a result:
@@
Finally, truncation may lead to an upward or downward biased CD index if the summation of equation (1) is negative and the inequalities of either equation (2) or equations (2) and (3) hold. In this case:

### v2

# Is There a Secular Decline in Disruptive Patents? Correcting for Measurement Bias

@@
Is There a Secular Decline in Disruptive Patents? Correcting for Measurement Bias
Jeffrey T. Macher Corresponding author. Robert E. McDonough School of Business, Georgetown University, 37th and O Streets NW, Washington, DC 20057, USA; jeffrey.macher@georgetown.edu. Christian Rutzer Faculty of Business and Economics, University of Basel, Peter Merian-Weg 6, 4002 Basel, Switzerland; christian.rutzer@unibas.ch. and Rolf Weder Faculty of Business and Economics, University of Basel, Peter Merian-Weg 6, 4002 Basel, Switzerland; rolf.weder@unibas.ch.
(26. Februar 2026)
Zusammenfassung
Despite tremendous growth in the volume of new scientific and technological knowledge, the popular press has recently raised concerns that disruptive innovation is slowing. These dire prognoses were driven in part by Park et al. (2023), a Nature publication that uses decades of data and millions of observations coupled with a novel quantitative metric (the CD index) that characterizes innovation in science and technology as either consolidating or disruptive. We challenge the Park et al. (2023) patent findings, principally around concerns of truncation bias and exclusion bias. We show that 88 percent of the decrease in the average CD index over 1980-2010 reported by the authors can be explained by their truncation of all backward patent citations before 1976. We also show that this truncation bias varies by technology class. We further account for a change in U.S. patent law that allows for citations to patent applications in addition to patent grants—something ignored by the authors in their analysis—and update the analysis to 2016. We show that the number of highly disruptive patents has increased since 1980—particularly since 2008. Our results suggest caution in using the Park et al. (2023) patent findings and conclusions as a basis for research and decision-making in public policy, industry restructuring or firm reorganization aimed at altering the current innovation landscape.
@@ 1 Introduction
Continued innovation in science and technology is considered a bedrock for and driving force of growth and prosperity in most economies. A co-authored paper by Park, Leahey and Funk entitled, “Papers and Patents are Becoming Less Disruptive Over Time,” was recently published in Nature (2023). Given its title and findings, as well as the topic examined, the paper attracted significant and global media attention. Park et al. (2023) has been featured in hundreds of international newspapers and magazines Altmetric (2023). As examples, The Economist (2023) emphasized in a report on the changing nature of science that "Papers and Patents are Becoming Less Disruptive"; The New York Times (2023) asked "What Happened to All of Science’s Big Breakthroughs?"; and the Financial Times (2023) noted SScience is Losing its Ability to Disrupt".
@@
One reason for this attention is that the Park et al. (2023) results show marked declines in disruptive research and innovation over time—a finding complementary to research productivity concerns expressed by other authors Chu and Evans (2021); Bloom et al. (2020); Boeing and Hünermund (2020). [^1] Bloom et al. (2020) conclude in their own analysis that ”[t]aking the US aggregate number as representative, research productivity falls in half every 13 years: ideas are getting harder and harder to find”(p. 1138). Park et al. (2023) emphasize in their abstract that the "results suggest that slowing rates of disruption may reflect a fundamental shift in the nature of science and technology"(p. 138), and note in their conclusion that "this trend is unlikely to be driven by changes in citation practices or the quality of published work. Rather, the decline represents a substantive shift in science and technology, one that reinforces concerns about slowing innovative activity. We attribute this trend in part to scientists’ and inventors’ reliance on a narrower set of existing knowledge"(p. 142).
A second reason for this attention is that the Park et al. (2023) results suggest implications for the organization of the entire science and technology innovation process—from government research labs and universities to private and public enterprises. Park et al. (2023) even provide suggestions as to how universities and federal agencies can ßtrongly reward research qualityänd help scholars to ïnoculate themselves from the publish or perish culture, and produce truly consequential work"(pp. 143-144), and conclude that "[u]nderstanding the decline in disruptive science and technology more fully permits a much-needed rethinking of strategies for organizing the production of science and technology in the future"(p. 144).
@@
The authors’ findings unsurprisingly piqued the interests of researchers, commentators, and journalists on a global scale. Some offered potential explanations, such as increased pressures to apply for large (interdisciplinary) projects, rising administrative burdens, declining basic research funding, increasing risk-aversion, and heightened publishing pressures Rust (2023); Yanai and Lercher (2023). Other researchers raised concerns around the CD index used by Park et al. (2023) as an appropriate measure of disruptive innovation Petersen et al. (2023); Leibel and Bornmann (2023); Ruan et al. (2021); Bornmann et al. (2020). An important question arises, however, as to whether the Park et al. (2023) findings are accurate.
This paper does not challenge the CD index as a measure of disruptiveness of newly created knowledge. The CD index was developed by Funk and Owen-Smith (2017) and can broadly be understood as a measure of whether a patent (scientific paper) is more Consolidating or Disruptive, based on the characteristics of its citations. A patent (scientific paper) is disruptive if "the subsequent work that cites it is less likely to also cite its predecessors"(Park et al., 2023, p. 139). Our concerns instead arise from how the authors truncate cited references (i.e., "backward citations") to patents and exclude citations to patent applications. We thus challenge their methodological approach due to truncation bias and exclusion bias.
@@
Restricting our analysis to the same patent data, we find that the worrisome result suggested by Park et al. (2023) is mainly a consequence of the omission of citations to older innovations: the authors truncate all backward patent citations before 1976. We also show that substantive differences in the results occur from not considering patent law changes: the authors exclude all patent application citations published after 2000. Both of these biases have large measurement effects on the CD index and direct implications on the authors’ findings and conclusions.
Correcting for truncation bias substantially decreases the apparent secular decline in patent disruptiveness suggested by Park et al. (2023): i.e., our analysis indicates a marginal decrease in the CD index over time. Correcting further for exclusion bias challenges the "remarkable stability"(p. 140) in the number of highly-disruptive patents argument suggested byPark et al. (2023): i.e., our analysis indicates an increase in the number of highly disruptive patents over time—particularly since 2008.222This finding differs from Kalyani (2022) who uses the appearance of new ”bigramsäs an indicator of novel innovation ideas, but is consistent with Braguinsky et al. (2023) who observe a sharp increase in ”novelÜnited States Patent and Trademark Office (USPTO) patent applications filed since 2005.
The rest of this paper is organized as follows. Section 2 compares the Park et al. (2023) methodology (with truncation of backward citations) to our methodology (without truncation of backward citations) using the same patent data as Park et al. (2023). Section 3 illustrates why truncation bias is concerning, and Section 4 assesses its significance. Section 5 updates the data and analysis over 2011-2016 and considers a 1999 patent law change that allows for citations to patent grants and to patent applications to assesses the significance of exclusion bias. Section 6 shows that the number of highly disruptive patents has increased over time—particularly since 2008. Section 7 offers concluding remarks and suggests caution against any proposed research policy changes in response to Park et al. (2023).
@@ 2 Measuring Patent Disruptiveness
The CD index was developed by Funk and Owen-Smith (2017) and used by Park et al. (2023) to characterize whether a patent or scientific publication is considered more consolidating (i.e., building upon previous research and reinforcing the status quo) or more disruptive (i.e., obsolescing previous research and pushing into new directions). [^3] Since its introduction by Funk and Owen-Smith (2017) the CD index has been used in a wide variety of analyses to capture the disruptive content of patents (e.g., Lin et al., 2023; Kaltenberg et al., 2023; Wu et al., 2019) and of scientific papers (e.g., Lin et al., 2023; Bornmann et al., 2020; Wu et al., 2019). The CD index ranges from -1 (consolidating) to 1 (disruptive). Park et al. (2023) use five-year post-publication windows in its construction, referred to as $CD_{5}$, and their analysis starts in 1945 for scientific papers and in 1980 for patents.
Fig. 1 plots two average $CD_{5}$ indices for approximately $3.66$ million USPTO patents over 19801980-$2010$: one index replicates the Park et al. (2023) methodology; the other index uses our methodology. [^4] As in Park et al. (2023), we consider only utility patents and draw from PatentsView (version February $21$, $2023$). This database contains all US patents published between $1976$ and September $29$, $2022$. While Park et al. (2023) focus on patents of the aggregate NBER technology fields ”Chemical”, ”Computer and Communications”, ”Pharmaceutical and Medical”, Ëlectrical and Electronic”, and ”Mechanical”that encompasses $3,046,672$ granted utility patents over 1980-2010, our main analysis utilizes all granted USPTO utility patents published that encompasses $3,662,051$ granted utility patents over the same period. See Appendix A for more details. As is readily apparent, the evolution of each index differs markedly: the Park et al. (2023) methodology produces an average annual CD index that starts at 0.39 in 1980 (the first observation year) but declines rapidly to 0.05 in 2010; our methodology instead produces an average annual CD index that starts at 0.09 in 1980 and declines more gradually to 0.05 in 2010. Fig. 1 nonetheless shows marked convergence in the two indices in the later sample years—particularly since 2000.
We express changes in these respective CD indices by using percentiles across their full range: i.e., -1 (0th percentile), 0 (50th percentile), and 1 (100th percentile) as shown on the right vertical axis of Fig. 1. The Park et al. (2023) methodology indicates the CD index declines from the 69.5 percentile in 1980 to the 52.5 percentile in 2010: a substantive decrease of 17 percentage points relative to its maximum possible change. [^5] The percentile of a CD index value can be calculated by 100*(CD index/2 + 0.5). Our methodology indicates the CD index declines from the 54.5 percentile in 1980 to the 52.5 percentile in 2010: a marginal decrease of two percentage points relative to its maximum possible change.
Abbildung 1: The CD index: Park et al. (2023) Methodology Versus Our Methodology
@@
The contrast in these CD indices would be even greater had Park et al. (2023) fully exploited the patent data available to them over 1976-2017. The authors indicate that they end their analysis before 2017 "because some measures require data from subsequent years for calculation"(p. 145): i.e., a five-year post-publication window. [^6] This five-year post-publication window does allow, in principle, for an average $CD_{5}$ index up to 2012; the authors nevertheless end their analysis in 2010. For illustration, suppose all of the available patent data were used up to 1976 instead of starting at 1980. The dotted line in the gray-shaded area in Fig. 1 using the Park et al. (2023) methodology indicates a CD index that declines from the 99.5 percentile in 1976 to the 52.5 percentile in 2010 (i.e., 47 percentage points), whereas our methodology indicates a CD index that declines from the 56.0 percentile to the 52.5 percentile over the same period (i.e., 3.5 percentage points).
The differences observed in the average CD indices over time are due to Park et al. (2023)’s truncation of all backward citations to patents published before 1976. Consequently, the results presented by Park et al. (2023) are biased by this truncation, with the bias strongest for those patents published closer to the truncation year. Their findings of a sharp decline in the average patent disruptiveness can mostly be attributed to a reduction in measurement bias; not to a decline in disruptive innovations. We find that 88 percent of the total decline in the average disruptiveness of patents over 1980-2010 found by Park et al. (2023) can be explained by the authors’ artificial truncation of all backward citations to patents published before 1976.777The percentage decrease due to truncation is calculated as follows: (1) the $CD_{5}$ values using the Park et al. (2023) Methodology are 0.39 in 1980 and 0.05 in 2010 – a 0.34 difference; (2) the $CD_{5}$ values using our No Truncation Methodology are 0.09 in 1980 and 0.05 in 2010 – a 0.04 difference. The difference in differences between the methods is 0.34 - 0.04 or 0.3. Expressed as a percentage: 0.3 / 0.34 * 100 = 88% of the total decrease in average disruptiveness.
@@ 3 Why Truncation Matters
An example helps to illustrate our argument: We compare two patents that both appear in the Park et al. (2023) dataset: [^8] Appendix Fig. 3(b) shows the first page from each patent document listing the backward citations. US 4181011 was issued in 1980; US 6511791 was issued in 2003. US 4181011 makes 13 citations to patents published between 1958 and 1974; US 6511791 makes 11 citations to patents published between 1987 and 2001. Using the Park et al. (2023) methodology and replication data (patentsview_analytical_df.csvpatentsview\_analytical\_df.csv), the number of backward citations for patent US 4181011 is zero—given the exclusion of backward citations to patents published before 1976. In contrast, the number of backward citations for patent US 6511791 is 11—the same number as listed in the patent document.
Abbildung 2: How Truncation of Backward Citations Affects the CD Index
@@
This can be shown formally as follows. Park et al. (2023) define the CD index for a focal patent as:
$CD_{t}=\frac{1}{N}\sum_{i=1}^{N}(-2f_{it}b_{it}+f_{it}),\penalty 10000\ with$ (1)
[…]
@@
[…]
where $N$ is the sum of the number of patents citing: the focal patent only ($N_{F}$), the focal patent and parts of its predecessors ($N_{B})$, and parts of its predecessors only ($N_{R}$) within $t$ years after the focal patent publication. A forward citation increases the consolidating (disruptive) nature of the focal patent if the forward cited patent does (does not) cite the predecessor patents (i.e., those patents also cited by the focal patent). [^9] Wu et al. (2019) and subsequent work by Bornmann et al. (2020) simplified the CD index formula as: $CD=\frac{N_{F}-N_{B}}{N_{F}+N_{B}+N_{R}}$.
We extend the CD index by directly considering truncation, which affects the CD index in two ways. First, left truncation reduces the number of counted forward citations $N$ if the truncated backward citations are cited within the $t$ post-publication years of the focal patent:
$N^{tr}<N^{non-tr}$ (2)
where $tr$ represents truncated and nontrnon-tr represents non-truncated backward citations.
Second, left truncation reduces the $b_{it}$ value if a forward citation of the truncated predecessor is also a forward citation of the focal patent within $t$ years of publication. In this case:
@@
Using equations (2) and (3), we consider four distinct cases: (i) no backward citations are truncated; (ii) backward citations are truncated but none are cited within the $t$ post-publication years of the focal patent; (iii) backward citations are truncated and cited within the $t$ post-publication years of the focal patent but none cite the focal patent; and (iv) backward citations are truncated and cited within the $t$ post-publication years of the focal patent and at least one of them cite the focal patent. In case (i) and case (ii), equations (2) and (3) hold as equalities and truncation has no effect on the CD index:
@@
In case (iii), the inequality of equation (2) holds but equation (3) is strictly an equality. In case (iv), the inequalities of equations (2) and (3) hold. In case (iii) and case (iv), the truncated CD index is biased upward as long as the summation part of equation (1) of the truncated CD index is positive, because each of the two effects artificially increases its value. Moreover, the bias is stronger in case (iv), ceteris partibus. Both results are easily seen by adding the inequalities of equations (2) and (3) to the CD index provided in equation (1). As a result:
@@
Truncation may therefore lead to upward or downward bias in the CD index if the summation of equation (1) is negative and the inequalities of either equation (2) or equations (2) and (3) hold. In this case: